5m+n =10
4m + n = 4

Answer by ikleyn(52797)

.
5m + n = 10
4m + n =  4 
----------------------


Distract the second equation from the first one. You will get

m = 6.

Hence, n = 10-5*m = 10-5*6 = 10-30 = -20.